OpenStudy (anonymous):

write in lowest terms : 4-x / 2x-8

OpenStudy (anonymous):

\[\frac{ 4-x }{ 2x-8 }\]

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

Not quite.

terenzreignz (terenzreignz):

\[\large \frac{4-x}{2x-8}=\frac{-(x-4)}{2(x-4)}\]

OpenStudy (anonymous):

@terenzreignz is right.

terenzreignz (terenzreignz):

Missing signs... all math students fall to that trap... myself included ^.^